OpenStudy (anonymous):

Find a ordered pair for... 5x+y=6

OpenStudy (anonymous):

@csscgirl

OpenStudy (anonymous):

where is the y in the equation? or is there a y in the equation?

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

its all good :D

OpenStudy (anonymous):

i fixed it

OpenStudy (anonymous):

found three... (0,6),(1,1),(2,-4)

OpenStudy (anonymous):

5x+7=6 I will try to solve your equation. 5x+7=6(Simplify both sides of the equation) 5x+7+−7=6+−7(Add -7 to both sides) 5x=−1 (5x) 5 = (−1) 5 (Divide both sides by 5) x=−0.2

OpenStudy (anonymous):

@sudonick rockon bro

OpenStudy (anonymous):

will any of those 3 work? @sudonick

OpenStudy (anonymous):

@csscgirl tyvm :)

OpenStudy (anonymous):

yea they should

OpenStudy (anonymous):

yup they work :D

OpenStudy (anonymous):

Okay thanks!